OpenStudy (anonymous):

If cosA+sinA=√2cosA,prove that cosA-sinA=√2sinA

OpenStudy (shubhamsrg):

cosA + sinA = √2 cosA sinA = √2 cosA - cosA sinA = cosA (√2 - 1) sinA/cosA = √2 - 1 tanA = √2 - 1 cotA = 1/(√2 - 1) cotA = 1/(√2 - 1) * (√2 + 1)/(√2 + 1) cotA = (√2 + 1)/(2 - 1) cotA = √2 + 1 cosA/sinA = √2 + 1 cosA = sinA (√2 + 1) cosA = √2 sinA + sinA cosA - sinA = √2 sinA I hope this helps!